Session: Nov 2003 (session 3)
Wavelength: 3.6 cm


To give feedback on this experiment, click here

Experiment Feedback

Medicina (Vargiu at Tue Oct 28 17:05:41 2003) Success :

Onsala (Michael Lindqvist at Tue Oct 28 17:10:48 2003) Success :
No known problems. Mark5+4 in piggyback.

Hartebeesthoek (Jonathan Quick at Tue Oct 28 17:25:28 2003) Success :
No known problems

Shanghai (Xinyong Huang at Tue Oct 28 22:26:09 2003) Success :
We lost five scans ( 2 scans antenna with troubel and 3 scans recorder with troubel).

Urumqi (yangwj at Wed Oct 29 02:13:59 2003) Minor failures : PARI
In the tape2, parity have some high. The signal of x-band have some instable,the reason is not clear,still need to check.

Effelsberg (A. Kraus at Wed Oct 29 08:23:50 2003) Minor failures : RECO
The first two scans have been lost due to Mark5-trouble. Between UT2200-2240, computer problems (not related to the VLBI system) led to the loss of data.

Noto (P. Cassaro at Mon Nov 3 13:15:01 2003) Minor failures : RECO
The recorder lost the vacuum. From 14:15 to 15:40 the data are lacking.

(Other) (Mario Berube at Mon Nov 3 16:29:07 2003) Success :
Algonquin completed EC017c without any known problems. All sources observed.

Ny_Alesund (Leif Morten Tangen at Wed Nov 5 07:37:52 2003) Success :
A lot of off-source alarms because of strong wind the first 12 hours.

MPIfR_Bonn (Walter Alef at Mon Nov 17 13:09:37 2003) Minor failures :
Noto amplitude low by a factor of about 14 on the baseline in S-band and by about 100 in X-band. Pointing Problem? Focus offset? Shanghai no fringes, no phase-cal.